OpenStudy (ashking1):

Solve the proportion 35/28-x/4 . What is the value of x

OpenStudy (anonymous):

5

OpenStudy (anonymous):

for this one, you want to just simplify the fraction on the left.factorize it and then eliminate the factors that are the same in the numerator and denominator

jhonyy9 (jhonyy9):

cross multiplie so than 4*35=28x so than x=4*35/28 =35/7 =5 so x=5

OpenStudy (ashking1):

thks

jhonyy9 (jhonyy9):

yw good luck bye